# A new approach to $P-V$ phase transitions: Einstein gravity and holographic dark energy

### v1

@@ III Conclusions
The aim of this work was to demonstrate the existence of $P-V$ phase transitions within the thermodynamics description of Einstein gravity for the case of a flat FLRW spacetime; as far as we know, this is the first work in which this result is established. We focused our attention on late times cosmic evolution, therefore the content of the universe is described by two dominant fluids: dark matter and dark energy. In our scheme, the fluids of the dark sector are not allowed to interact but are related throughout the usual cosmological coincidence parameter, which we denote by $r$. Since the parameter $r$ is, in general, expected to be a dynamical function, we must be able to write it in terms of the geometry of the apparent horizon, i.e., it must be a function of the radius, $R_{A}$, which is the dynamical quantity for an expanding cosmology. Additionally to the consideration of two fluids, this is a compelling requirement for the parameter $r$ in order to guarantee the existence of $P-V$ phase transitions in our approach. Furthermore, the prerequisite mentioned above for $r$ is backed by the conditions that establish the existence of critical points in the equation of state of the universe, namely $P=P(V,T)$; such conditions are defined by means of partial derivatives w.r.t. $R_{A}$.
@@
According to our construction, the form of the dark energy density is restricted to be proportional to a power-law function of the Hubble parameter, $H$. We restrict ourselves to values of the exponent $n$ that are physically justified. For instance, $n=1$ represents the well-known holographic dark energy widely explored in the literature with characteristic length given by the Hubble scale. The case $n=2$ is a dark energy model that leads to an early phantom scenario in Einstein’s gravity, and the cases $n=3,4$ resemble the correction terms appearing in the acceleration equation of the cosmic evolution emerging from modified gravity theories in 4D and those in which cubic and quartic invariants of the curvature tensor are taken into account. Therefore, this simple form of the dark energy density is able to encompass several cases of interest. We also extended our analysis to the case in which a linear term modifies the dark energy density in the first derivative of $H$. This kind of contribution is properly studied within the holographic dark energy scenario (Ricci and Granda-Oliveros models). In QCD, this type of modification appeared in the context of the so-called ghost dark energy.
@@
For $n=1$, no critical points are found; consequently, no $P-V$ phase transitions are observed. We attribute this result to the cosmological coincidence parameter exhibiting a constant behavior. In this case, the $P(V,T)$ equation of state does not differ from the usual expression discussed in Einstein’s gravity. However, for $n>1$, the isotherms displayed in the $P-V$ diagram show the coexistence phase for temperature values below the critical temperature, and the isotherms over the critical temperature indicate an ideal gas behavior as the temperature increases. Therefore, a $P-V$ phase transition is expected for the cosmological fluid, and according to the Gibbs free energy behavior, we have a first-order phase transition. On the other hand, as the value of the exponent $n$ increases we observe that the thermodynamics instability is attenuated, in other words, the size of the loop characterizing the first-order transition decreases, this means that in the $P-V$ diagram the isotherms that show coexistence phase are smoothed. Therefore, higher powers in $H$ could lead to scenarios with no phase transitions, and this characteristic provides a cutoff in the value of the exponent $n$. An interesting result is that the thermodynamics description carried out does not depend on the value of the arbitrary parameter introduced in the dark energy density, the variations obtained between different cases are due to the value of the exponent $n$ only.
@@ Acknowledgements.
M. Cruz work has been supported by S.N.I.I. (CONAHCyT-México). J. Saavedra acknowledges the financial support of Fondecyt Grant 1220065. M. Cruz also acknowledges the warm hospitality of the Physics Institute of PUCV and valuable discussions with F. Tello-Ortiz during the preparation of this manuscript.

### v2

@@ III Conclusions
The aim of this work was to demonstrate the existence of $P-V$ phase transitions within the thermodynamic description of Einstein gravity for the case of a flat FLRW spacetime; as far as we know, this is the first work in which this result is established. We focused our attention on cosmic evolution in the late times. Therefore, the universe’s content is described by two dominant fluids: dark matter and dark energy. In our scheme, the fluids of the dark sector are not allowed to interact but are related throughout the usual cosmological coincidence parameter, which we denote by $r$. Since the parameter $r$ is, in general, expected to be a dynamical function, we must be able to write it in terms of the geometry of the apparent horizon, i.e., it must be a function of the radius, $R_{A}$, which is the dynamical quantity for an expanding cosmology. Considering two fluids, this is a compelling requirement for the parameter $r$ to guarantee the existence of $P-V$ phase transitions in our approach. Furthermore, the prerequisite mentioned above for $r$ is backed by the conditions that establish the existence of critical points in the equation of state of the universe, namely $P=P(V,T)$; such conditions are defined through partial derivatives w.r.t. $R_{A}$.
@@
According to our construction, the form of the dark energy density is restricted to be proportional to a power-law function of the Hubble parameter, $H$. We restrict ourselves to values of the exponent $n$ that are physically justified. For instance, $n=1$ represents the well-known holographic dark energy widely explored in the literature with characteristic length given by the Hubble scale. The case $n=2$ is a dark energy model that leads to an early phantom scenario in Einstein’s gravity, and the cases $n=3,4$ resemble the correction terms appearing in the acceleration equation of the cosmic evolution emerging from modified gravity theories in 4D and those in which cubic and quartic invariants of the curvature tensor are taken into account. However, other holographic models can be easily adapted to our analysis, see for instance hdemodels where the Barrow, Tsallis, Renyi, and Sharma-Mittal (to mention some) are dark energy models which fulfill the condition, $\rho_{\mathrm{de}}\propto H$. Therefore, this simple assumption on the form of the dark energy density can encompass several cases of interest. We also extended our analysis to the case in which a linear term modifies the dark energy density in the first derivative of $H$. This contribution is properly studied within the holographic dark energy scenario (Ricci and Granda-Oliveros models). In QCD, this type of modification appeared in the context of the so-called ghost dark energy.
@@
For $n=1$, no critical points are found; consequently, no $P-V$ phase transitions are observed. We attribute this result to the cosmological coincidence parameter exhibiting a constant behavior. In this case, the $P(V,T)$ equation of state does not differ from the usual expression discussed in Einstein’s gravity. However, for $n>1$, the isotherms displayed in the $P-V$ diagram show the coexistence phase for temperature values below the critical temperature, and the isotherms over the critical temperature indicate an ideal gas behavior as the temperature increases. Therefore, a $P-V$ phase transition is expected for the cosmological fluid, and according to the Gibbs free energy behavior, we have a first-order phase transition. On the other hand, as the value of the exponent $n$ increases we observe that the thermodynamic instability is attenuated, in other words, the size of the loop characterizing the first-order transition decreases, this means that in the $P-V$ diagram the isotherms that show coexistence phase are smoothed. Therefore, higher powers in $H$ could lead to scenarios with no phase transitions, and this characteristic provides a cutoff in the value of the exponent $n$. An interesting result is that the thermodynamic description does not depend on the value of the arbitrary parameter introduced in the dark energy density. The variations obtained between different cases are due to the value of the exponent $n$ only.
@@
Finally, we emphasize that there are still a lot of subtle issues worth investigating when tests against observations are implemented for the holographic dark energy models treated in this work, for instance, the restriction of the exponent $n$ appearing in (25) with astrophysical data. However, we feel that any further exploration along this line would justify a separate analysis.
@@ Acknowledgements.
M. Cruz work has been supported by S.N.I.I. (CONAHCyT-México). J. Saavedra acknowledges the financial support of Fondecyt Grant 1220065. M. Cruz also acknowledges the warm hospitality of the Physics Institute of PUCV and valuable discussions with F. Tello-Ortiz during the preparation of this manuscript. Anonymous reviewers are acknowledged for helpful suggestions.
Appendix A Classical stability
Following the line of reasoning of Ref. stability , we study the stability of the background evolution of the dark energy models arising from Eq. (25) for $n=1,2,3,4$ by analyzing the sign of the squared speed of sound given by
$v^{2}=\frac{dp_{\mathrm{de}}}{d\rho_{\mathrm{de}}},$ (43)
where $v^{2}<0$ represents classical instability of the model at perturbation level. For a barotropic equation of state the above equation is simply
$v^{2}=\omega_{\mathrm{de}}+\rho_{\mathrm{de}}\frac{d\omega_{\mathrm{de}}}{d\rho_{\mathrm{de}}}.$ (44)
Inserting the dark energy (25) in the continuity equation $\dot{\rho}_{\mathrm{de}}+3H\rho_{\mathrm{de}}(1+\omega_{\mathrm{de}})=0$, one gets
$\omega_{\mathrm{de}}=-1+\frac{2n}{3}(1+q)=-1+2n\left(\frac{\rho_{\mathrm{m}}}{2\rho_{\mathrm{m}}-(2n-1)\rho_{\mathrm{de}}}\right),$ (45)
where we have used the Friedmann constraint (17) to write the second equality. Therefore we can write Eq. (44) as follows
$v^{2}=\frac{4n}{3-2n}\left(\frac{r^{2}(z)}{2r(z)-2n+1}\right),$ (46)
being $r(z)$ the coincidence parameter defined previously. Notice that for the epoch of dark energy dominance the coincidence parameter is restricted to the interval $(0,1]$ where $r\approx 0$ at the future ($z<0$) and $r\approx 1$ at early times ($z<1$) as well as present time ($z=0$); thus for $n=1$ we have $v^{2}<0$ and for $n=2,3,4$ we obtain $v^{2}>0$ during early and present times cosmic evolution, respectively. According to our results the case $n=1$ it lacks of physical interest since no $P-v$ phase transition is obtained and in turn results to be classically unstable. Regarding the model given by Eq. (37) we follow the discussion given in Ref. unstable . In general, the consideration of the term $\dot{H}$ induces instabilities in the holographic dark energy model. However, the introduction of more parameters into the holographic energy density can lead to stable scenarios. We consider that in our case different values for the exponent $n$ could have the same effect as the extra parameter (Barrow exponent) considered and discussed in unstable .